# DualAnoDiff: Dual-Interrelated Diffusion Model for Few-Shot Anomaly Image Generation

### v1

@@ Abstract
The performance of anomaly inspection in industrial manufacturing is constrained by the scarcity of anomaly data. To overcome this challenge, researchers have started employing anomaly generation approaches to augment the anomaly dataset. However, existing anomaly generation methods suffer from limited diversity in the generated anomalies and struggle to achieve a seamless blending of this anomaly with the original image. In this paper, we overcome these challenges from a new perspective, simultaneously generating a pair of the overall image and the corresponding anomaly part. We propose DualAnoDiff, a novel diffusion-based few-shot anomaly image generation model, which can generate diverse and realistic anomaly images by using a dual-interrelated diffusion model, where one of them is employed to generate the whole image while the other one generates the anomaly part. Moreover, we extract background and shape information to mitigate the distortion and blurriness phenomenon in few-shot image generation. Extensive experiments demonstrate the superiority of our proposed model over state-of-the-art methods in terms of both realism and diversity. Overall, our approach significantly improves the performance of downstream anomaly detection tasks, including anomaly detection, anomaly localization, and anomaly classification tasks.
Introduction
Industrial anomaly inspection, i.e., anomaly detection, localization, and classification, plays an important role in industrial manufacture (Duan et al. 2023). However, in real-world industrial production, anomaly samples are scarce. Therefore, the current mainstream anomaly inspection methods are either unsupervised methods which use only normal samples (Roth et al. 2022; Li et al. 2024) or semi-supervised method (Zhang et al. 2023) which employ both the normal samples and a few anomaly data. Although these methods perform well in anomaly detection, they have limited performance in anomaly localization and can not deal with the task of anomaly classification (Hu et al. 2024). Therefore, researchers proposed anomaly generation methods to generate more anomaly data, to help achieve better performance by using supervised anomaly inspection.
@@ Introduction
Figure 1: Top: Evaluating anomaly generation quality in four aspects: Whether generate valid anomaly, degree of realism, alignment of the mask, and whether the location of mask is reasonable, the results show that our generated results are better than the other methods. Bottom: Our model can simultaneously generate extensive anomaly image-mask pairs.
Existing anomaly generation methods can be categorized into two groups, 1) model-free methods randomly cut and paste patches from existing anomalies or anomalous texture datasets onto normal samples (Lin et al. 2021; Li et al. 2021). But the anomaly data synthesized by them is unrealistic. 2) Generative methods employ generative models like GANs and diffusion models to generate anomaly data. Generally, The GAN-based model (Zhang et al. 2021; Niu et al. 2020) requires a large amount of training data to achieve better generation performance and they can not generate masks. DFMGAN (Duan et al. 2023) firstly trained on normal data and then migrated to anomaly data to achieve a few shot generation. This method also encounters the problem that the generated anomalies are not realistic enough and the masks are not sufficiently aligned because there is no explicit alignment constraint design. AnomalyDiffusion (Hu et al. 2024) based on the texture-inversion (Gal et al. 2022) technique of Diffusion (Rombach et al. 2022), separately learns the anomaly appearance and location information, then generates the anomaly on the masked normal samples. As AnomalyDiffusion only focuses on the part of anomaly, the generated anomalies do not blend realistically with the original image, and the masks generated individually may appear in the background of image.
@@
To address these limitations, we propose DualAnoDiff, a novel few-shot anomaly image generation model that utilizes a dual-diffusion to simultaneously generate the overall image and the corresponding anomaly part. This approach can realize the effective integration of anomaly image and anomaly part, resulting in the generation of authentic and highly aligned anomaly image-mask data pairs. Specifically, our model is built upon a pre-trained diffusion model and introduces two LoRA to expand a single diffusion model into dual-interrelated diffusion model. One branch, referred to as the global branch, is responsible for generating the overall anomaly image, while the anomalous branch generates the localized anomaly image. They exchange information through the self-attention interaction module, This module combines the attention layers of global and anomalous branches and performs shared attention calculations, enabling the interaction and fusion of information in the dual-denoising process. This ensures the consistency between the generated overall anomaly image and the localized anomaly image. Furthermore, to further preserve the invariance of the background, we introduce a background compensation module based on self-attention adaptive injection. This module involves adding noise to the background image, extracting the key and value from the intermediate feature layer, and applying adaptive fusion MLP to incorporate the background information into the global branch. It helps to maintain consistency in the generated image and avoids coupling between the anomaly part and the background of the object in the image.
@@
Fig. 1 shows our generation results outperform better compared with other method in four key aspects. Moreover, extensive experiments have been conducted in Mvtec AD (Bergmann et al. 2019) to quantitatively validate the superiority of the anomaly data generated by DualAnoDiff in downstream anomaly inspection tasks, and achieving a state-of-the-art level of performance with pixel-level 99.1% AUROC and 84.5% AP score.
@@
  • We design a background compensation approach, which involves image backgrounds as control information and injects the intermediate feature into the denoising process of anomaly image, to enhance the stability and authenticity of the generated data.
@@
Related Work
Few-shot Image Generation
@@
Few-shot image generation aims to generate new and diverse examples while preventing overfitting to the few training images (Zhu et al. 2024). It is highly susceptible to overfit with extremely limited training data (less than 10) and then generate highly similar images. Early method (Mo, Cho, and Shin 2020) proposes modifying network weights, using various regularization techniques and data augmentation to prevent overfitting (Hu et al. 2024). There are also methods (Hu et al. 2023) to mitigate overfitting by pre-training on the source domain and subsequently migrating to fewer samples through cross-domain consistency losses to keep the generated distribution. Textual Inversion (Gal et al. 2022) and Dreambooth (Ruiz et al. 2023) encode a few images into the textual space of a pre-trained diffusion model to achieve diverse target customization generation which preserves its key visual features.
Anomaly Inspection
@@
The anomaly inspection task consists of anomaly detection, localization, and classification (Duan et al. 2023). Due to the scarcity of abnormal data in industrial scenarios, most methods (Roth et al. 2022; Li et al. 2024) use unsupervised methods and semi-supervised methods. Reconstruction-based methods (Schlegl et al. 2019; Cao et al. 2023) detect anomalies by analyzing the residual image before and after reconstruction. Embedding-based methods (Lee, Lee, and Song 2022; Cao et al. 2022) utilize pre-trained networks to extract the image-level features and patch-level features, and then perform clustering according to the similarity between the features to detect the anomalies. All of those methods can only address the task of anomaly detection, while having limited performance in anomaly localization and being incapable of anomaly classification.
Anomaly Generation
@@
Due to the scarcity of anomaly data, anomaly generation has emerged as a field of paramount importance. DRAEM (Zavrtanik, Kristan, and Skočaj 2021), Cut-Paste (Li et al. 2021), Crop-Paste (Lin et al. 2021) and PRN (Zhang et al. 2023) crop and paste unrelated textures or existing anomalies into normal sample. These approaches can be somewhat effective, but the generated anomalies are completely unrealistic and there is no way to use them as anomaly classification tasks. Subsequently, GANs (Goodfellow et al. 2020) have been applied for anomaly generation due to their ability to generate high-fidelity images. SDGAN (Niu et al. 2020) and DefectGAN (Zhang et al. 2021) generate anomalies on normal samples by learning from anomaly data. However, they require a large amount of anomaly data and cannot generate anomaly masks. DFMGAN (Duan et al. 2023) transfers a StyleGAN2 (Karras et al. 2020) pretained on normal samples to the anomaly domain, but lacks generation authenticity and accurate alignment between generated anomalies and masks. Subsequently, Diffusion models have been more widely used due to their extreme generalizability. AnomalyDiffusion (Hu et al. 2024) learns the features of anomaly and the distribution of masks through text inversion (Gal et al. 2022) technique of diffusion, to generate the specified anomaly at the position of the corresponding mask of the normal image. However, Since the method learns the anomaly part and the mask separately, it makes the generated mask not necessarily located on the object, and the anomaly-object transition is not natural enough. Our approach is similarly based on the diffusion model but achieves more realistic, distinct, and diverse anomaly data generation.
Method
@@
Preliminaries
Latent Diffusion Models. Stable Diffusion (SD), a variant of the latent diffusion model (LDM) (Rombach et al. 2022), serves as a text-guided diffusion model. To generate high-resolution images while enhancing computational efficiency in the training process, it employs a pre-trained variational autoencoder (VAE) (Kingma and Welling 2013) encoder $\mathcal{E}(\cdot)$ to map images into latent space and perform an iterative denoising process. Subsequently, the predicted images are mapped back into pixel space through the pre-trained VAE decoder $D(\cdot)$. $\epsilon_{\theta}$ is the denoising network, for each denoising step, the simplified optimization objective is defined as follows:
@@
where $\epsilon$ are latent noise, the text description $c$ is encoded by the CLIP (Radford et al. 2021) text encoder $\tau_{\theta}(\cdot)$ and then used to guide the diffusion denoising process.
DualAnoDiff Framework
We have been searching for methods that can generate both anomaly image and mask, inspired by Layerdiffusion (Zhang and Agrawala 2024), we decompose an anomaly image into two parts, the overall anomaly image and the corresponding anomaly part, where the overall image refers to the whole anomaly image $I$, the anomaly part refers to the part that contains only the anomaly $I_{a}$. $I_{a}=I\times M_{a}$ ($M_{a}$ is the mask of anomaly).
As shown in Fig. 2, the proposed DualAnoDiff involves two interrelated diffusion models, and they share part information by the Self-attention Interaction Module. We freeze the weights of diffusion models and use two LORA (Hu et al. 2021) to fine-tune them. For ease of description, we denote the two diffusion models as $SD$ and $SD^{\ast}$, $SD$ denote the diffusion model to generate image $I$, $SD^{\ast}$ denote the diffusion model to generate $I_{a}$.
Dual-Interrelated Diffusion. AnomalyDiffusion (Hu et al. 2024) primarily focuses on the anomaly part, which may result in generated anomaly images lacking a convincingly realistic appearance. Directly generating the complete anomaly image poses challenges in obtaining the corresponding mask. To address those limitations, our proposed model simultaneously generates both the overall image and the anomaly part. This novel approach overcomes challenges in generating realistic anomaly images while ensuring the availability of accurate masks.
Firstly, we encode $I$, $I_{a}$ into latent space $z$,$z^{\prime}$ = $\varepsilon(I),\varepsilon(I_{a})$ by using VAE encoder $\varepsilon(\cdot)$, employ a forward process to add noise into the latents with the same timestep $t$, and then learn to denoising during the backward process guided by different prompts. Throughout these processes, information is shared and synchronized between the two diffusion models through the SAIM, enabling the model to effectively fit the training data pairs. By generating the anomalies separately, this approach achieves two important objectives. Firstly, it enhances the diversity and reality of the generated anomaly through a simple yet effective operation of adding two LoRA. Secondly, it ensures a highly aligned mask that accurately corresponds to the anomaly image.
Nested Prompts. The goal of the dual diffusion is to generate an anomaly image and anomaly part pair ($I$ and $I^{\prime}$), which exhibits an inclusion relationship. To facilitate the model’s understanding of the distinct entities within the image (the primary subject and the anomaly), we employ a pair of prompts designed to reflect this inclusion relationship:
@@
where the prompts $p$ and $p^{\prime}$ correspond to the anomaly image $I$ and the corresponding anomaly part $I_{a}$ respectively. Both prompts are encoded by the text encoder $\tau_{\theta}(\cdot)$ and then injected into Unet (Ronneberger, Fischer, and Brox 2015) of Diffusion. $\tau_{\theta}(\cdot)$ is trainable.
@@
Figure 3: b, c are the cross attention maps of different text tokens in SD corresponding to the text of “a vfx with” and “sks”. a, d are the generated image from SD and SD*.
@@
The variables $x$ and $y$ can be the class name and anomaly name provided by the dataset. In our model, we use the $vfx$ and $sks$ which were suggested by DreamBooth (Ruiz et al. 2023). Those words have weak prior in both the language model and the diffusion model, making them easier to fit than other words, and can achieve better generation results, specially for high-prior words. Fig. 3 presents the generated result and visualizes the cross-attention maps between text token and vision. where $a$, $d$ are the anomaly image and anomaly part generated by $SD$ and $SD^{\ast}$. $b$ and $c$ are the 64 $\times$ 64 resolutions feature maps randomly extracted from the second half of the generation process in the U-Net of $SD$. Where $b$ corresponds to the text “$a$ $vfx$ $with$”, $c$ corresponds to “$sks$”. It is evident that the model correctly separates the attributes of anomaly and object, and accurately associating them to the specified text as we want.
bottle 1.52 0.04 1.43 0.04 1.57 0.06 1.39 0.07 1.62 0.12 1.58 0.19 2.17 0.36
cable 1.97 0.19 1.74 0.25 1.89 0.19 1.70 0.22 1.96 0.25 2.13 0.41 2.12 0.43
capsule 1.37 0.06 1.23 0.05 1.49 0.03 1.59 0.04 1.59 0.11 1.59 0.21 1.60 0.31
grid 1.97 0.07 2.00 0.12 1.95 0.10 2.01 0.12 1.97 0.13 2.04 0.44 2.09 0.42
hazel_nut 1.97 0.05 1.74 0.21 1.85 0.16 1.87 0.19 1.93 0.24 2.13 0.31 1.91 0.35
leather 1.80 0.07 1.47 0.14 2.04 0.12 2.12 0.14 2.06 0.17 1.94 0.41 1.88 0.34
metal_nut 1.55 0.04 1.56 0.15 1.45 0.28 1.47 0.30 1.49 0.32 1.96 0.30 1.56 0.32
pill 1.56 0.06 1.49 0.11 1.61 0.07 1.61 0.10 1.63 0.16 1.61 0.26 1.82 0.37
screw 1.13 0.11 1.12 0.16 1.17 0.10 1.19 0.12 1.12 0.14 1.28 0.30 1.34 0.36
tile 2.10 0.12 1.83 0.20 2.53 0.21 2.35 0.22 2.39 0.22 2.54 0.55 2.35 0.50
transistor 1.61 0.13 1.39 0.15 1.76 0.13 1.47 0.13 1.64 0.25 1.57 0.34 1.69 0.33
wood 2.05 0.03 1.95 0.23 2.12 0.25 2.19 0.29 2.12 0.35 2.33 0.37 2.21 0.40
zipper 1.30 0.05 1.23 0.11 1.25 0.10 1.25 0.10 1.29 $\underline{0.27}$ 1.39 0.25 2.09 0.36
Average 1.65 0.07 1.51 0.14 1.71 0.13 1.69 0.15 1.72 0.20 1.80 0.32 1.90 0.37
Self-attention Interaction Module (SAIM). During training, $SD$ and $SD^{\ast}$ share the same timestep $t$ and denoising simultaneously, they share information after every attention in the unet. For example, self-attention is more likely to share the positional information and cross-attention shares the semantic information. The shared step is formulated as:
$\varphi_{i}(\tilde{z})=$ $Rearrange(Cancat(\varphi_{i}(z),\varphi_{i}(z^{\prime})))$
$\varphi_{i}(\tilde{z})_{new}=$ $SelfAtt(\varphi_{i}(\tilde{z}))$
$\varphi_{i}(z),\varphi_{i}(z^{\prime})=$ $Split(Rearrange(\varphi_{i}(\tilde{z})_{new}+\varphi_{i}(\tilde{z})))$ (3)
@@
where $\varphi^{i}$ is the intermediate representation of the U-Net. The original shape of $\varphi_{i}(z)$ and $\varphi_{i}(z^{\prime})$ are “b w c”, where b denotes the batch size, w represents the spatial dimension, and c represents the channel dimension. We rearrange the $cancat(\varphi_{i}(z),\varphi_{i}(z^{\prime}))$ to shape of “bw 2 c”. Concat and split are a set of corresponding operations that are used to aggregate and separate $varphi_{i}(z),\varphi_{i}(z^{\prime}))$.
@@
Figure 4: Comparison between the models without (a) and with (b) the Background Compensation Module
@@
Loss Function. With the two-stream parallel structure of diffusion, our final training objective expressed as follows:
@@
here, $\epsilon$ and $\epsilon^{\ast}$ are latent noise for the anomaly image and anomaly part, $t$ is the time step, $p$, $p^{\prime}$ are the corresponding prompts, they are encoded by text encoder $\tau_{\theta}(\cdot)$ which is trainable.
Mask Generation. Thanks to our two-stream parallel structure, which generates the anomaly part as a single entity, obtaining the precise mask becomes straightforward. There are two ways to obtain a high-quality mask. Firstly, we can utilize existing segmentation algorithms such as SAM (Kirillov et al. 2023), U2-Net (Qin et al. 2020) to obtain high-quality masks after generating the anomaly part image. Additionally, during the generation process, we can extract the average attention maps in $SD^{\ast}$ to compute the mask such as (Wang et al. 2023) which is widely used in semantic segmentation. In our paper, we use the first method.
Background Compensation Module
Although the model has achieved good results so far, there are still some problems in some cases, which is caused by the limited training data. Several bad cases of corresponding generated results are shown in Fig. 4 for categories bottle, pill, and toothbrush in the Mvtec (the average training data for each category is 8). The bottle was only partially generated. The edges of pill is blurred and the internal properties leakage into the background. Additionally, the toothbrush has two brush handles, which is an abnormal occurrence. Furthermore, all of those generated images lack sharpness and the background color does not maintain sufficient purity. In general, the model is incapable of fully grasping the characteristics of those cases, there are problems such as objects and background mixed, and objects deformed.
To enhance the model’s ability to learn the shape of object and focus more on object generation, we design a background compensation module. First, We employ U2-Net to segment the image $I$ and obtain the object mask $M_{f}$, then get the background image $I_{b}=(1-M_{f})\times I$, $I_{b}$ will process the $SD$ as same as $I$ (except for the SAIM). Next, We utilize $I_{b}$ as a condition, which contains both background and mask information. This allows us to to control the shape of the object while providing background context. We collect its features in self-attentions and inject them into the $K$, $V$ of the $I$ in every corresponding self-attention steps. The injection process can be formulated as:

### v3

# Dual-Interrelated Diffusion Model for Few-Shot Anomaly Image Generation

@@ Abstract
The performance of anomaly inspection in industrial manufacturing is constrained by the scarcity of anomaly data. To overcome this challenge, researchers have started employing anomaly generation approaches to augment the anomaly dataset. However, existing anomaly generation methods suffer from limited diversity in the generated anomalies and struggle to achieve a seamless blending of this anomaly with the original image. Moreover, the generated mask is usually not aligned with the generated anomaly. In this paper, we overcome these challenges from a new perspective, simultaneously generating a pair of the overall image and the corresponding anomaly part. We propose DualAnoDiff, a novel diffusion-based few-shot anomaly image generation model, which can generate diverse and realistic anomaly images by using a dual-interrelated diffusion model, where one of them is employed to generate the whole image while the other one generates the anomaly part. Moreover, we extract background and shape information to mitigate the distortion and blurriness phenomenon in few-shot image generation. Extensive experiments demonstrate the superiority of our proposed model over state-of-the-art methods in terms of diversity, realism and the accuracy of mask. Overall, our approach significantly improves the performance of downstream anomaly inspection tasks, including anomaly detection, anomaly localization, and anomaly classification tasks. Code will be made available.
1 Introduction
@@
Figure 1: Top: Evaluating anomaly generation quality in four aspects: Whether generate valid anomaly, degree of realism, alignment of the mask, and whether the location of mask is reasonable, the results show that our generated results are better than the other methods. (Yellow area represents the valid generation, green indicates the correct mask or area, red indicates the generated mask or wrong area.) Bottom: Our model can simultaneously generate extensive anomaly image-mask pairs.
Industrial anomaly inspection, i.e., anomaly detection, localization, and classification, plays an important role in industrial manufacture [6]. However, in real-world industrial production, anomaly samples are scarce. Therefore, the current mainstream anomaly inspection methods are either unsupervised methods which use only normal samples [36, 22] or semi-supervised method [48] which employ both the normal samples and a few anomaly data. Although these methods perform well in anomaly detection, they have limited performance in anomaly localization and can not deal with the task of anomaly classification [14]. Therefore, researchers proposed anomaly generation methods to generate more anomaly data, to help achieve better performance by using supervised anomaly inspection.
@@
Existing anomaly generation methods can be categorized into two groups, 1) model-free methods randomly cut and paste patches from existing anomalies or anomalous texture datasets onto normal samples [23, 20, 46]. But the anomaly data synthesized by them are unrealistic. 2) Generative methods employ generative models like GANs and diffusion models to generate anomaly data. Generally, GAN-based model [47, 26] requires a large amount of training data to achieve better generation performance and they can not generate masks. DFMGAN [6] is firstly trained on normal data and then migrated to anomaly data to achieve a few-shot generation. This method also encounters the problem that the generated anomalies are not realistic enough [29] and the masks are not sufficiently aligned [21] because there is no explicit alignment constraint design. AnomalyDiffusion [14] based on the texture-inversion [7] technique of Diffusion [33], separately learns the anomaly appearance and location information, then generates the anomaly on the masked normal samples. As AnomalyDiffusion only focuses on the part of anomaly, the generated anomalies do not blend realistically with the original image, and the masks generated individually may appear in the background of image. To address these limitations, we propose DualAnoDiff, a novel few-shot anomaly image generation model that utilizes a dual-interrelated diffusion to simultaneously generate the overall image and the corresponding anomaly part. This approach can realize the effective integration of anomaly image and anomaly part, resulting in the generation of realistic and highly aligned anomaly image-mask data pairs with good diversity. Specifically, our model is built upon a pre-trained diffusion model and introduces two LoRA [12] to expand a single diffusion model into dual-interrelated diffusion model. One branch, referred to as the global branch, is responsible for generating the overall anomaly image, while the anomalous branch generates the localized anomaly image. They exchange information through the self-attention interaction module. This module combines the attention layers of global and anomalous branches and performs shared attention calculations, enabling the interaction and fusion of information in the dual-denoising process. This ensures the consistency between the generated overall anomaly image and the localized anomaly image. Furthermore, to further preserve the invariance of the background, we introduce a background compensation module based on self-attention adaptive injection. This module involves adding noise to the background image, extracting the key and value from the intermediate feature layer, and applying adaptive fusion MLP to incorporate the background information into the global branch. It contributes to maintaining the accuracy of the background and the shape of the object in the generated images, while avoiding the coupling between the object and the background in the images.
@@
Fig.7 shows our generated results outperform those of other methods in four key aspects. Moreover, extensive experiments have been conducted on MVTec AD [1] to quantitatively validate the superiority of the anomaly data generated by DualAnoDiff in downstream anomaly inspection tasks, and achieving a state-of-the-art level of performance in pixel-level anomaly detection with 99.1% AUROC and 84.5% AP score.
@@
  • We design a background compensation approach, which involves image backgrounds as control information and injects the intermediate feature into the denoising process of anomaly image, to enhance the stability and realism of the generated data.
@@
2 Related Work
2.1 Few-shot Image Generation
@@
Few-shot image generation aims to generate new and diverse examples while preventing overfitting to the few training images [51, 45]. It is highly susceptible to overfit with extremely limited training data (less than 10) and then generate highly similar images. FreezeD [25] proposes modifying network weights, using various regularization techniques and data augmentation to prevent overfitting [14]. There are also methods [13] to mitigate overfitting by pre-training on the source domain and subsequently migrating to fewer samples through cross-domain consistency losses to keep the generated distribution. Textual Inversion [7] and Dreambooth [37] encode a few images into the textual space of a pre-trained diffusion model to achieve diverse target customization generation which preserves its key visual features. Although these methods can generate realistic images, they are incapable of generating pixel-level annotations which are essential for anomaly image generation tasks. In contrast, our method enables high-quality annotations to be readily obtained by simultaneously generating local anomaly images.
2.2 Anomaly Inspection
@@
The anomaly inspection task consists of anomaly detection, localization, and classification [6]. Due to the scarcity of abnormal data in industrial scenarios, most methods [36, 9, 22, 18, 24, 42, 40, 41] use unsupervised methods and semi-supervised methods. Reconstruction-based methods [39, 3, 10] detect anomalies by analyzing the residual image before and after reconstruction. Embedding-based methods [19, 2, 44] utilize pre-trained networks to extract the image-level features and patch-level features, and then perform clustering according to the similarity between the features to detect the anomalies. All of those methods can only address the task of anomaly detection, while having limited performance in anomaly localization and being incapable of anomaly classification. Through the generation of abnormal images, these three tasks can be successfully accomplished, and our method achieves the state-of-the-art performance.
2.3 Anomaly Generation
@@
Due to the scarcity of anomaly data, anomaly generation has emerged as a field of crucial significance. DRAEM [46], Cut-Paste [20], Crop-Paste [23] and PRN [48] crop and paste unrelated textures or existing anomalies into normal sample. These approaches can be somewhat effective, but the generated anomalies are completely unrealistic and there is no way to use them as anomaly classification tasks. Subsequently, GANs [8] have been applied for anomaly generation due to their ability to generate high-fidelity images. SDGAN [26] and DefectGAN [47] generate anomalies on normal samples by learning from anomaly data. However, they require a large amount of anomaly data and cannot generate anomaly masks. DFMGAN [6] transfers a StyleGAN2 [15] pretained on normal samples to the anomaly domain, but lacks generation realism and accurate alignment between generated anomalies and masks. Subsequently, Diffusion models have been more widely used due to their extreme generalizability. AnomalyDiffusion [14] learns the features of anomaly and the distribution of masks through text inversion [7] technique of diffusion, to generate the specified anomaly at the position of the corresponding mask of the normal image. However, Since the method learns the anomaly part and the mask separately, it makes the generated mask not necessarily located on the object, and the anomaly-object transition is not natural enough. Nevertheless, through the utilization of a two-branch diffusion model along with a background supplement module, our methodology successfully accomplishes the full decoupling of diverse attributes within the anomaly image. As a result, the generated data are more realistic and diversified, and the performance of various downstream tasks is significantly enhanced.
3 Method
@@
3.1 Preliminaries
Latent Diffusion Models. Stable Diffusion (SD), a variant of the latent diffusion model (LDM) [33], serves as a text-guided diffusion model. To generate high-resolution images while enhancing computational efficiency in the training process, it employs a pre-trained variational autoencoder (VAE) [16] encoder $\mathcal{E}(\cdot)$ to map images into latent space and perform an iterative denoising process. Subsequently, the predicted images are mapped back into pixel space through the pre-trained VAE decoder $D(\cdot)$. $\epsilon_{\theta}$ is the denoising network, for each denoising step, the simplified optimization objective is defined as follows:
@@ 3.1 Preliminaries
where $\epsilon$ are latent noise, the text description $c$ is encoded by the CLIP [31] text encoder $\tau_{\theta}(\cdot)$ and then used to guide the diffusion denoising process.
3.2 DualAnoDiff Framework
We have been searching for methods that can generate both anomaly image and mask. Inspired by Layerdiffusion [49], we decompose an anomaly image into two parts, the overall anomaly image and the corresponding anomaly part, where the overall image refers to the whole anomaly image $I$, the anomaly part refers to the part that contains only the anomaly region $I_{a}$. $I_{a}=I\times M_{a}$ ($M_{a}$ is the mask of anomaly part).
As shown in Fig. 2, the proposed DualAnoDiff involves two interrelated diffusion models, and they share part information by the Self-attention Interaction Module. We freeze the weights of diffusion models and use two LORA [12] to fine-tune them. For ease of description, we denote the two diffusion models as $SD$ and $SD^{\ast}$, $SD$ denote the diffusion model to generate global image $I$, $SD^{\ast}$ denote the diffusion model to generate anomaly part $I_{a}$.
Dual-Interrelated Diffusion. AnomalyDiffusion [14] primarily focuses on the anomaly part, which may result in generated anomaly images lacking a convincingly realistic appearance. However, generating the complete anomaly image poses challenges in obtaining the corresponding mask. To address those limitations, our proposed model simultaneously generates both the overall image and the anomaly part. This novel approach overcomes challenges in generating realistic anomaly images while ensuring the availability of accurate masks.
First, we encode $I$, $I_{a}$ into latent space $z$ and $z^{\prime}$, with $z=\varepsilon(I)$ and $z^{\prime}=\varepsilon(I_{a})$ by using VAE encoder $\varepsilon(\cdot)$. Next we employ a forward process to add noise into the latents with the same timestep $t$, and then learn to denoising during the backward process guided by different prompts. Throughout these processes, information is shared and synchronized between the two diffusion models through the SAIM, enabling the model to effectively fit the training data pairs. By generating the anomalies separately, this approach achieves two important objectives. Through a simple yet effective operation of adding two LoRA, it enhances the diversity and reality of the generated anomaly. Additionally, it ensures a highly aligned mask that accurately corresponds to the anomaly image.
Nested Prompts. The goal of the dual diffusion is to generate an anomaly image and anomaly part pair ($I$ and $I_{a}$), which exhibits an inclusion relationship. To facilitate the model’s understanding of the distinct entities within the image (the primary subject and the anomaly), we employ a pair of prompts designed to reflect this inclusion relationship:
@@ 3.2 DualAnoDiff Framework
where the prompts $p$ and $p^{\prime}$ correspond to the anomaly image $I$ and the corresponding anomaly part $I_{a}$ respectively. Both prompts are encoded by the trainable text encoder $\tau_{\theta}(\cdot)$ and then injected into Unet [34] of Diffusion.
@@
Figure 3: a is the image generated by SD, b and c are the cross attention maps of different text tokens in SD corresponding to the text of “a vfx with” and “sks”.
@@
The variables $x$ and $y$ can be the class name and anomaly name provided by the dataset. In our model, we use the $vfx$ and $sks$ which were suggested by DreamBooth [37]. Those words have weak prior in both the language model and the diffusion model, making them easier to fit than other words, and can achieve better generation results, specially for high-prior words. Fig.3 presents the generated result and visualizes the cross-attention maps between text token and vision. where $a$ is the anomaly image generated by $SD$. $b$ and $c$ are the 64 $\times$ 64 resolutions feature maps randomly extracted from the second half of the generation process in the Unet of $SD$. Where $b$ corresponds to the text “$a$ $vfx$ $with$”, $c$ corresponds to “$sks$”. It is evident that the model correctly separates the attributes of anomaly and object, and accurately associating them to the specified text as we want. Self-attention Interaction Module (SAIM). During training, $SD$ and $SD^{\ast}$ share the same timestep $t$ and denoising simultaneously, they share information by SAIM after every attention blocks in the Unet of diffusion. For example, after the self-attention blocks is more likely to share the positional information and detailed information, the cross-attention blocks shares the semantic information.
In SAIM, we use attention to fuse the information from two branch, The shared step is formulated as:
$\varphi_{i}(\tilde{z})=Rearrange$ $(Concat(\varphi_{i}(z),\varphi_{i}(z^{\prime})))$
$\varphi_{i}(\tilde{z})_{new}=$ $SelfAtt(\varphi_{i}(\tilde{z}))$ (3)
$\varphi_{i}(z)^{\prime},\varphi_{i}(z^{\prime})^{\prime}=Split($ $Rearrange(\varphi_{i}(\tilde{z})_{new}+\varphi_{i}(\tilde{z})))$
@@
where $\varphi_{i}$ is the intermediate representation of the Unet. The original shape of $\varphi_{i}(z)$ and $\varphi_{i}(z^{\prime})$ are “b w c”, where “b” denotes the batch size, “w” represents the spatial dimension, and “c” represents the channel dimension. We rearrange the $Concat(\varphi_{i}(z),\varphi_{i}(z^{\prime}))$ to shape of “bw 2 c” to avoid displacement in space. $Concat$ and $Split$ are a set of corresponding operations that are used to aggregate and separate feature maps.
@@
Figure 4: Comparison between the models without (a) and with (b) the Background Compensation Module.
@@
Loss Function. With the two-stream structure of diffusion, our final training objective is expressed as follows:
@@
where, $\epsilon$ and $\epsilon^{\ast}$ are latent noise for the anomaly image and anomaly part, $t$ is the time step, $p$, $p^{\prime}$ are the corresponding prompts, they are encoded by text encoder $\tau_{\theta}(\cdot)$ which is trainable.
Mask Generation. Thanks to our two-stream parallel structure, which generates the anomaly part as a single entity, obtaining the precise mask becomes straightforward. There are two ways to obtain a high-quality mask: 1). We utilize existing segmentation algorithms such as SAM [17], U2-Net [30] to obtain high-quality masks after generating the anomaly part image. 2) During the generation process, we can extract the average attention maps in $SD^{\ast}$ to compute the mask such as [43] which is widely used in semantic segmentation. In this paper, we use the first method.
3.3 Background Compensation Module
Although the model has achieved good results so far, there are still some problems in some cases, which is caused by the limited training data. Several bad cases of corresponding generated results are shown in Fig.4 for categories bottle, pill, and toothbrush in the MVTec (the average training data for each category is 8). The bottle was only partially generated. The edges of pill is blurred and the internal properties leakage into the background. Additionally, the toothbrush has two brush handles, which is an abnormal occurrence. Furthermore, all of those generated images lack sharpness and the background color does not maintain sufficient purity. In general, since the model is incapable of fully grasping the characteristics of those cases, there are problems such as objects being mixed with the background and objects being deformed.
To enhance the model’s ability to learn the shape of object and focus more on object generation, we design a background compensation module. First, we employ U2-Net to segment the image $I$ and obtain the object mask $M_{f}$, then get the background image $I_{b}=(1-M_{f})\times I$, $I_{b}$ will be processed by the $SD$ as same as $I$ (except for the SAIM). Then, we utilize $I_{b}$ as a condition, which contains both background and mask information. This allows us to to control the shape of the object while providing background context. We collect its features in self-attentions and inject them into the $K$, $V$ of the $I$ in every corresponding self-attention steps. The injection process can be formulated as: